OpenStudy (anonymous):

A dormitory manager buys 38 bed sheets and 61 towels for $791.50. The manager get another 54 bed sheets and 50 towels for $923 from the same store. What is the cost of one bed sheet and one towel? $15 and $10 $13.50 and $7.50 $12 and $5.50 $10.50 and $3

OpenStudy (abmon98):

Purchases- 38 bed Sheets 61 Towels Price of purchases- $791.50 Purchases- 54 bed Sheets 50 Towels Price of purchases- $923.00 38x+61y=791.50 54x+50y=923.00 y=791.50-38x/61 54x+50(791.50-38x/61)=923 54x+(39575-1900x/61)=923 3294x+39575-1900x=56303 3294x-1900x=16728 1394x=16728 x=12.00

OpenStudy (abmon98):

y=791.50-(38*12)/61 y=5.50